Exam 6: Probability

arrow
  • Select Tags
search iconSearch Question
flashcardsStudy Flashcards
  • Select Tags

If P(A)= .30,P(B)= .60,and P(A and B)= .20,then P(A|B)= .40.

(True/False)
4.8/5
(77)

Suppose A and B are two independent events for which P(A)= 0.20 and P(B)= 0.60. a.Find P(A and B). b.Find P(A or B).

(Short Answer)
4.7/5
(90)

Two events A and B are said to be mutually exclusive if:

(Multiple Choice)
4.8/5
(74)

We can apply the multiplication rule to compute the probability that two events occur at the same time.

(True/False)
4.8/5
(82)

____________________ can find the probability that someone with a disease tests positive by using (among other things)the probability that someone who actually has the disease tests positive for it.

(Short Answer)
4.8/5
(71)

In the scenario of Bayes' Law,P(A|B)is a posterior probability,while P(B|A)is a(n)____________________ probability.

(Short Answer)
4.9/5
(77)

When A and B are mutually exclusive,P(A or B)can be found by adding P(A)and P(B).

(True/False)
4.8/5
(93)

If the two events are mutually exclusive and collectively exhaustive,what is the probability that one or the other occurs?

(Multiple Choice)
4.8/5
(81)

The probability of the intersection is called a joint probability.

(True/False)
4.8/5
(82)

GPA and Class A college professor classifies his students according to their grade point average (GPA)and their class rank.GPA is on a 0.0-4.0 scale,and class rank is defined as the under class (freshmen and sophomores)and the upper class (juniors and seniors).One student is selected at random.  GPA and Class  A college professor classifies his students according to their grade point average (GPA)and their class rank.GPA is on a 0.0-4.0 scale,and class rank is defined as the under class (freshmen and sophomores)and the upper class (juniors and seniors).One student is selected at random.   ​ ​ -{GPA and Class Narrative} Are being in the upper class and having a GPA over 3.0 related? Explain. ​ ​ -{GPA and Class Narrative} Are being in the upper class and having a GPA over 3.0 related? Explain.

(Essay)
4.7/5
(91)

Hobby Shop Sales Sales records of a hobby shop showed the following number of radio controlled trucks sold weekly for each of the last 50 weeks.  Hobby Shop Sales Sales records of a hobby shop showed the following number of radio controlled trucks sold weekly for each of the last 50 weeks.   ​ ​ -{Hobby Shop Sales Narrative} What is the probability of selling at least two trucks in any given week? ​ ​ -{Hobby Shop Sales Narrative} What is the probability of selling at least two trucks in any given week?

(Short Answer)
4.8/5
(76)

Is it possible to have two events for which P(A)= 0.40,P(B)= 0.50,and P(A or B)= 0.30? Explain.

(Essay)
4.9/5
(72)

We can use the joint and marginal probabilities to compute conditional probabilities,for which a formula is available.

(True/False)
4.8/5
(63)

The ____________________ rule says that P(Ac)= 1 − P(A).

(Short Answer)
4.8/5
(69)

Marital Status An insurance company has collected the following data on the gender and marital status of 570 customers. Marital Status An insurance company has collected the following data on the gender and marital status of 570 customers.   suppose that a customer is selected at random. -{Marital Status Narrative} Are gender and marital status mutually exclusive? Explain using probabilities. suppose that a customer is selected at random. -{Marital Status Narrative} Are gender and marital status mutually exclusive? Explain using probabilities.

(Essay)
4.8/5
(68)

In general,a posterior probability is calculated by adding the prior and likelihood probabilities.

(True/False)
4.8/5
(66)

Drunk Drivers Six hundred accidents that occurred on a Saturday night were analyzed.Two items noted were the number of vehicles involved and whether alcohol played a role in the accident.The numbers are shown below: Drunk Drivers  Six hundred accidents that occurred on a Saturday night were analyzed.Two items noted were the number of vehicles involved and whether alcohol played a role in the accident.The numbers are shown below:   ​ ​ -{Drunk Drivers Narrative} What proportion of accidents involved more than one vehicle? ​ ​ -{Drunk Drivers Narrative} What proportion of accidents involved more than one vehicle?

(Short Answer)
4.8/5
(77)

Messenger Service ​ Three messenger services deliver to a small town in Oregon.Service A has 60% of all the scheduled deliveries,service B has 30%,and service C has the remaining 10%.Their on-time rates are 80%,60%,and 40% respectively.Define event O as a service delivers a package on time. ​ ​ -{Messenger Service Narrative} If a package was delivered on time,what is the probability that it was service A?

(Essay)
4.8/5
(88)

Two events A and B are said to be mutually exclusive if P(A and B)= 1.0.

(True/False)
4.7/5
(77)

Two events A and B are said to be independent if P(A|B)= P(B).

(True/False)
5.0/5
(72)
Showing 61 - 80 of 240
close modal

Filters

  • Essay(0)
  • Multiple Choice(0)
  • Short Answer(0)
  • True False(0)
  • Matching(0)